OpenStudy (anonymous):

9,98,987,9876.... compleet this Consecutive and find Un=8*.......?

OpenStudy (anonymous):

ok its hard to find it look at this u1=8*1+1=9 u2=8*12+2=98 u3=8*123+3=987 u4=8*1234+4=9876 .. ..